Calculate the mass of 0.105mol Na2CO3.10H2O


1
Expert's answer
2022-01-24T13:15:05-0500

Solution:

The molar mass of Na2CO3·10H2O is 286.14 g/mol

Hence,

Mass of Na2CO3·10H2O = (0.105 mol) × (286.14 g / 1 mol) = 30.045 g

Mass of Na2CO3·10H2O = 30.045 g


Answer: 30.045 g of Na2CO3·10H2O